### v2

# Function-Correcting Codes for Insertion-Deletion Channels

Abstract
In coding theory, handling errors that occur when symbols are inserted or deleted from a transmitted message is a long-standing challenge. Optimising redundancy for insertion and deletion channels remains a key open problem, with significant importance for applications in DNA data storage and document exchange. Recently, a new coding framework known as function-correcting codes has been proposed to address the challenge of optimising redundancy while preserving particular functions of the message. This framework has gained attention due to its potential applications in machine learning systems and long-term archival data storage. To address the problem of redundancy optimisation in insertion and deletion channels, we propose a new coding framework called function-correcting codes for insertion-deletion channels.
In this paper, we introduce the notions of function-correcting insertion codes, function-correcting deletion codes, and function-correcting insdel (insertion–deletion) codes(FCIDCs), and we demonstrate that these three formulations are equivalent. We then introduce insdel distance matrices and irregular insdel-distance codes, and further derive lower and upper bounds on the optimal redundancy achievable by function-correcting codes for insdel channels. Furthermore, we establish Gilbert-Varshamov and Plotkin-like bounds on the length of irregular insdel-distance codes. By utilising the relation between optimal redundancy and the length of irregular insdel-distance codes, we provide another simplified lower bound on optimal redundancy. We subsequently find bounds on optimal redundancy of FCIDCs for various classes of functions, including locally bounded functions, VT syndrome functions, the number-of-runs function, and the maximum-run-length function.
Keywords: Function-correcting codes, error-correcting codes, insertion-deletion channels, optimal redundancy.
1 Introduction
Traditional error-correcting codes(ECCs) are designed to enable the decoder to recover the transmitted message exactly. However, in many real-world situations, the decoder simply needs to compute a certain function of the message rather than reconstructing the complete message. The authors of [1] were inspired by this realization to develop a new class of codes called function-correcting codes (FCCs), which encode the message so that the decoder can reliably compute the desired function value, even in the presence of errors, with substantially less redundancy than required by classical ECCs for full message recovery.
FCCs have been commonly studied under systematic encoding, which is important in applications such as distributed computing and long-term archival storage, where preserving the original data is essential. Since redundancy is the primary and most crucial factor for FCCs, it becomes crucial to precisely determine optimal redundancy—that is, the minimum number of additional bits required to guarantee correct computation of the intended function. To determine optimal redundancy for both generic and specific functions, the authors of [1] introduced the concept of irregular-distance codes, which are closely connected to the redundancy patterns of FCCs. They established equivalence between the optimal redundancy of FCCs and the shortest achievable length of irregular-distance codes. Using this equivalence, the authors derived redundancy bounds for arbitrary functions and subsequently used these results to obtain tight or near-tight bounds for several important function classes, including locally binary bounded functions, Hamming weight functions, Hamming weight distribution functions, and real-valued functions.
1.1 Motivation
The nature of the underlying channel noise fundamentally determines how function values are distorted during transmission. By designing FCCs that are tailored to specific channel models, one can achieve reliable function computation in scenarios where recovering the entire message is either unnecessary or prohibitively costly in redundancy. Moreover, such channel-specific constructions provide deeper insight into the way different error processes affect function evaluation. Motivated by this observation, we focus on selecting suitable channels and associated functions to construct function-correcting codes with improved redundancy. In particular, we consider insertion–deletion(insdel) channels, where errors occur when symbols are inserted into or deleted from a transmitted message. Such errors disrupt the alignment between transmitted and received sequences, creating a form of noise that traditional error-correcting codes, designed primarily for handling substitution errors, are not well-suited to handle. As these types of errors frequently arise in biological systems, modern storage devices, and asynchronous communication settings, the study of insdel channels has become increasingly important in developing robust and efficient coding techniques. The study of insertion–deletion errors began in the 1960s [2, 3, 4], and [5], when V. Levenshtein introduced the edit distance, defining the minimum number of insertions, deletions, or substitutions required to transform one string into another. This concept established the mathematical basis for analyzing synchronization errors, although early work focused mainly on error detection and correction rather than determining full channel capacities. Later, practical coding schemes for handling such errors began to emerge, most notably the Varshamov–Tenengolts (VT) codes, originally introduced in [2] and subsequently expanded, enabling reliable single-deletion correction. During this period, researchers also explored the use of synchronizing sequences to help maintain alignment between transmitted and received data, marking a significant step toward systematically managing insertion-deletion errors. Consequently, the development of function-correcting codes under the insdel metric represents both a substantial theoretical milestone and a vital practical capability for next-generation communication and computation systems. These codes extend the reach of conventional error-control techniques beyond the constraints of the Hamming metric by providing resilience against synchronization errors, particularly insertions and deletions. This enables more reliable information processing, transmission, and storage in developing high-density, high-throughput environments. The nature of the underlying noise fundamentally influences how function values are distorted during transmission. A function that is robust to substitution errors may behave unpredictably under insertions, deletions, erasures, or symbol-dependent noise, and the structural interactions between the function and the channel can vary dramatically. Developing FCCs tailored to specific channels enables reliable computation in settings where full message recovery is either unnecessary or prohibitively expensive, while also revealing deeper insights into how functions transform under diverse error processes. Such channel-aware FCC frameworks are essential for modern applications ranging from asynchronous communication and DNA storage to distributed sensing and edge computing, where computation must remain accurate despite highly nonuniform and context-dependent noise.
1.2 Related Works
Function-correcting codes (FCCs) were introduced by Lenz et al. in [1] as a coding framework for recovering the value of a target function from the data rather than the data itself, resulting in substantially lower redundancy than traditional error-correcting codes. Their work formalized FCCs, established an equivalence with irregular-distance codes, and provided general upper and lower bounds demonstrating how the structure of the protected function influences the required redundancy. Subsequent developments, such as [6], revisited the model and extended the framework from the Hamming metric to the symbol-pair metric, deriving several bounds and structural insights. In [7], the authors established a lower bound on the redundancy of FCCs for linear functions. Focusing on the Hamming weight and Hamming weight distribution, Ge et al. [8] produced stronger lower bounds on redundancy and introduced code designs that exactly meet those bounds.
A growing line of research has further generalized FCCs to nonstandard error models. In particular, [9] extended function-correcting codes to the $b$-symbol read channel, addressing clustered read errors common in modern storage systems. That work introduced irregular $b$-symbol distance codes and derived bounds characterizing the redundancy needed to recover a function value under this metric. The results, supported by a graph-theoretic formulation and illustrative examples, show that FCCs can achieve substantially lower redundancy than traditional $b$-symbol error-correcting codes when only the function output must be protected.
Further extensions were provided in [10], which studied function-correcting codes for locally bounded $b$-symbol functions over $b$-symbol read channels. Recently, function correction has also been explored in the Lee-metric setting [11, 10]. More recent work, such as [12], investigates hybrid protection models in which both the function value and selected parts of the data must remain resilient to errors, highlighting the versatility of FCCs as a targeted and efficient alternative to full-message protection.
1.3 Contributions
Theoretical work on function-correcting codes has traditionally focused on channels such as the Hamming, symbol-pair, and Lee channels, all of which model substitution-type errors: a received symbol may differ from the transmitted one, but the positions of the symbols remain aligned.
The insertion–deletion (insdel) channel is fundamentally different because it introduces synchronization errors. Instead of altering symbol values, the channel may insert or delete symbols, shifting the entire sequence and breaking positional alignment between sender and receiver. As a result, a single insertion or deletion can misalign all subsequent symbols. Errors are measured under edit distance, not Hamming-type metrics. Decoding is considerably more difficult, since the code must recover both the correct symbols and the correct alignment.
These properties make designing and analyzing function-correcting codes for the insdel channel significantly more challenging than for traditional substitution-error channels. We develop the theoretical foundations of function-correcting codes for the insertion–deletion (insdel) channel and derive both lower and upper bounds on the optimal redundancy that can be achieved by such codes. Furthermore, we establish Gilbert–Varshamov-type and Plotkin-like bounds on the lengths of irregular insdel-distance codes. By leveraging the relationship between optimal redundancy and the lengths of irregular insdel-distance codes, we also provide a simplified lower bound on optimal redundancy.
Building on these results, we construct explicit Function-Correcting Insertion–Deletion Codes (FCIDCs) for several classes of functions, including locally bounded functions, VT functions, the number-of-runs function, and the maximum-run-length function.
1.4 Organization
The rest of the paper is organized as follows. Section II provides a comprehensive review of the foundational concepts used throughout the paper. We begin by recalling the basic definitions and properties of insertion–deletion (insdel) codes, including the insdel metric and several equivalent formulations that are relevant to our analysis. In Section III, we introduce the framework of function-correcting deletion codes, function-correcting insertion codes, and function-correcting insertion–deletion (insdel) codes, which are designed to recover the value of a target function in the presence of deletion, insertion, or combined insertion–deletion errors. We then present and discuss irregular insdel-distance codes, highlighting their significance in characterizing optimal redundancy and in deriving key performance limits. Building on this framework, we derive lower and upper bounds on the optimal redundancy achievable by function-correcting codes for insdel channels. In addition, we establish Gilbert–Varshamov-type and Plotkin-like bounds on the lengths of irregular insdel-distance codes, which play a central role in analysing and understanding the redundancy of such codes. In Section IV, we apply the general theoretical results to a specific class of functions, the VT syndrome function, which plays a central role in the insertion–deletion coding. We analyze how the structure of the VT syndrome function interacts with insdel errors and derive both upper and lower bounds on the optimal redundancy required for function correction in this setting. We illustrate the general framework using three concrete classes of functions: the number-of-runs function (Section V), the maximum-run-length function (Section VI), and locally bounded functions (Section VII). For each class, we derive the corresponding redundancy values and their associated upper and lower bounds. Section VIII concludes the paper.
1.4.1 Notations
Throughout this paper, we use the following notation:
  • $\mathbb{F}_{2}$: the binary field.
  • $\mathbb{F}_{2}^{n}$: the vector space of all binary vectors of length $n$.
  • $\mathbb{N}$: the set of positive integers; $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$: the set of non-negative integers.
  • For vectors $x,y$:
    • $\circ$
      $d_{H}(x,y)$: the Hamming distance,
    • $\circ$
      $d_{ID}(x,y)$: the insertion-deletion distance,
    • $\circ$
      $\operatorname{LCS}(x,y)$: the length of the longest common subsequence of $x$ and $y$.
  • $|x|$: the length of a vector (or string) $x$.
  • $D_{t}(x)$: the set of all subsequences obtained by deleting exactly $t$ symbols from $x$.
  • $I_{t}(x)$: the set of all supersequences obtained by inserting exactly $t$ symbols into $x$.
  • $r(x)$: the number of runs in a binary vector $x$.
  • Code parameters:
    • $\circ$
      $k$: message length,
    • $\circ$
      $r$: redundancy,
    • $\circ$
      $n=k+r$: total codeword length.
  • $\mathbb{N}_{0}^{M\times M}$: the set of all $M\times M$ matrices with non-negative integer entries.
  • $I]_{ij}$: the $(i,j)$-th entry of a matrix $I$.
  • For any integer $M$:
    • $\circ$
      $M]\triangleq\{1,2,\dots,M\}$,
    • $\circ$
      $M]^{+}\triangleq\max\{M,0\}$.
2 Preliminaries
In this section, we review the basic concepts and definitions used throughout the paper. We begin with a brief discussion of insertion, deletion, and insdel codes, which are designed to enable reliable communication over channels that produce synchronization errors. We then introduce the fundamental notions associated with irregular-distance codes, a class of codes characterized by non-uniform or structure-dependent distance measures. Throughout the paper, we focus on codes over the alphabet $\mathbb{F}_{2}=\{0,1\}$.
2.1 Insdel code
In this paper, we consider a channel model in which errors occur as deletions, insertions, or a combination of both. Codes designed to handle such types of errors were first introduced by Levenshtein, Varshamov, and Tenengolts in the 1960s [2, 3, 4], and [5]. In this section, we provide a brief overview of these classical codes and summarize the main results associated with them.
Let $x$ be a binary vector of length $k$ that, when transmitted through a deletion channel, may lose some of its bits. If the channel deletes $t$ bits, the resulting binary vector $y$ has length ktk-t. The vector $y$ is referred to as a subsequence of $x$. A subsequence of a vector is obtained by selecting a subset of its symbols and aligning them in their original order, without any rearrangement. Formally, the subsequence is defined as follows:
Definition 2.1 (Subsequence).
A sequence $x=x_{1}\ldots x_{k}$ is called subsequence of $y=y_{1}\ldots y_{n}$ if there are $k$ indices $i_{1}<\ldots<i_{k}$ such that $x_{1}=y_{i_{1}}\ldots x_{k}=y_{i_{k}}.$
Similarly, when $x$ is transmitted through an insertion channel, additional bits may be inserted at arbitrary positions, resulting in a supersequence of $x$. In the deletion case, the original vector $x$ is referred to as a supersequence of the received vector $y$. The set of all subsequence of $x$ obtained by deletion of $t$ bits is denoted by $D_{t}(x)$ and set of supersequences obtained by insertion of $t$ bits is denoted by $I_{t}(x)$.
We now define the main object of interest, which is deletion-correcting code and insertion-correcting code.
Definition 2.2 (Deletion Correcting Codes).
[3] A $t$-deletion-correcting code $\mathcal{C}$ of length $n$ is a subset of binary vector space $\mathbb{F}_{2}^{n}$ that holds the following property for all vectors $x,y\in\mathcal{C}$.
$D_{t}(x)\cap D_{t}(y)=\emptyset.$
Definition 2.3 (Insertion Correcting Codes).
[3] A $t$-insertion correcting code $\mathcal{C}$ of length $n$ is a subset of binary vector space $\mathbb{F}_{2}^{n}$ that holds the following property for all vectors $x,y\in\mathcal{C}$.
$I_{t}(x)\cap I_{t}(y)=\emptyset.$
Codes that are capable of correcting both insertions and deletions are referred to as insertion-deletion codes (or insdel codes). These codes can also be characterised through a suitable metric description, defined as follows:
Definition 2.4 (Insdel Metric).
[3] Let $x$ and $y$ be binary sequences of length $k$. The insdel metric $d_{ID}(x,y)$ between $x$ and $y$ is defined as the minimum number of insertions and deletions required to transform $x$ into $y$.
A sequence $z$ is a common subsequence of $x$ and $y$ if it is a subsequence of both $x$ and $y$. A longest common subsequence is a common subsequence of maximum possible length. Considering only insertion and deletion operations, we can define the insdel distance in terms of the length of the longest common subsequence as follows:
$d_{ID}(x,y)=|x|+|y|-2LCS(x,y)=2k-2LCS(x,y)$
where $|x|$ denotes the length of a binary word $x$ and $LCS(x,y)$ denotes the length of the longest common subsequence between $x$ and $y$.
Example 2.5.
Let $x=101$ and $y=010$ then the insdel distance between $x$ and $y$ is given as follows:
$101\xrightarrow[\text{at position 1}]{deletion}01\xrightarrow[\text{at position 3}]{insertion}010.$
Therefore $d_{ID}(x,y)=2$
Definition 2.6.
(Insdel Code)[13] A subset of $\mathbb{F}^{k}_{2}$ is defined to be an insdel code $\mathcal{C}$ whose minimum distance is given by:
$d_{ID}(\mathcal{C})=min_{c_{1},c_{2}\in\mathcal{C},c_{1}\neq c_{2}}{d_{ID}(c_{1},c_{2})}.$
The insdel distance of a code is a key parameter, as it determines the code’s capability to correct insdel errors. A code is said to be a $t$-insdel error-correcting code if its insdel distance is at least $2t+1$.
Remark 2.7.
[3] Any code that can correct $s$ deletions (or equivalently, any code that can correct $s$ insertions) can also correct $s$ combined deletions and insertions.
The following lemma gives a relation between the Hamming distance and the insdel distance.
Lemma 2.8.
[14] Let $x,y$ be two binary words of length $k$. If $d_{H}(x,y)$ denotes the hamming distance between $x$ and $y$ then we have:
$d_{ID}(x,y)\leq 2d_{H}(x,y).$
Insertion and deletion errors can disrupt the alternating structure of a binary string. A useful combinatorial invariant for studying such transformations is the run count. The following lemma gives a lower bound on the insdel distance in terms of the difference in run counts. This bound will later help us establish limits on code correction and code construction capabilities.
Lemma 2.9.
Let $x,y\in\mathbb{F}_{2}^{n}$ be two binary vectors of length $n$ and let $r(x)$ and $r(y)$ be the number of runs corresponding to vectors $x$ and $y$ respectively. Then,
$d_{ID}(x,y)\geq 2\Big\lceil\frac{|r(y)-r(x)|}{2}\Big\rceil.$
Moreover, when |r(x)r(y)||r(x)-r(y)| is odd then $d_{ID}(x,y)\geq|r(x)-r(y)|+1.$
Proof.
Assume without loss of generality that $r(y)\geq r(x)$. Let $m$ denote the number of insertions (and hence also the number of deletions) in an optimal insertion-deletion transformation from $x$ to $y$, so that the total number of operations is $d_{ID}(x,y)=2m$.
Let $A$ be the total change in run count caused by the $m$ insertions, and let $B$ be the total change in run count caused by the $m$ deletions. Then
r(y)r(x)=A+B.r(y)-r(x)=A+B.
Since a deletion can never increase the number of runs in a binary string, we have $B\leq 0$. Moreover, each insertion can increase the number of runs by at most $2$, and hence
$A\leq 2m.$
Combining these inequalities yields
$r(y)-r(x)=A+B\leq A\leq 2m.$
Therefore,
$m\geq\Big\lceil\frac{r(y)-r(x)}{2}\Big\rceil.$
Recalling that $d_{ID}(x,y)=2m$, we obtain
$d_{ID}(x,y)\geq 2\Big\lceil\frac{r(y)-r(x)}{2}\Big\rceil.$
If $r(x)\geq r(y)$, the same argument applies after swapping the roles of $x$ and $y$, yielding the bound in terms of $\lvert r(x)-r(y)\rvert$. Hence, in general,
$d_{ID}(x,y)\geq 2\Big\lceil\frac{\lvert r(x)-r(y)\rvert}{2}\Big\rceil.$
If $\lvert r(x)-r(y)\rvert$ is odd, then
$2\Big\lceil\frac{\Delta}{2}\Big\rceil=\Delta+1,$
where $\Delta=\lvert r(x)-r(y)\rvert$, completing the proof. ∎
We illustrate the tightness of the bound in Lemma 2.9 with two minimal examples, considering both even and odd run-count differences.
Example 2.10.
Even difference. Let $y=0100$ and $x=0000$. Here $r(y)=3$, $r(x)=1$, so |r(y)r(x)|=2|r(y)-r(x)|=2. The bound yields
$d_{ID}(x,y)\geq 2\Big\lceil\frac{2}{2}\Big\rceil=2.$
The insdel distance between $x$ and $y$ is given by
$d_{ID}(x,y)=2(4-3)=2.$
Hence, the bound is tight.
Odd difference. Let $y=0001$ and $x=0000$. Here $r(y)=2$, $r(x)=1$, so |r(y)r(x)|=1|r(y)-r(x)|=1. The bound yields
$d_{ID}(x,y)\geq 2\Big\lceil\frac{1}{2}\Big\rceil=2.$
The insdel distance between $x$ and $y$ is given by
$d_{ID}(x,y)=2(4-3)=2.$
Hence, the bound is tight in this case too.
For even number $2\leq d\leq 2k$, denote
$B_{ID}(x,d)\triangleq\{v\in\mathbb{F}_{2}^{k}:d_{ID}(x,v)\leq d\}$
as the insdel ball centered at $x$ of radius $d$. Then, we have the following bounds on $|B_{ID}(x,d)|$.
Lemma 2.11.
Let $k,t$ be positive integers $k\geq t\geq 1$. Then, for any $x\in\mathbb{F}_{2}^{k}$, it holds that:
$|B_{ID}(x,2t)|\leq\binom{r(x)+t-1}{t}\Bigg(\sum_{i=0}^{t}\binom{k}{i}\Bigg)\leq\Bigg(\frac{e^{2}(k+t)k}{t^{2}}\Bigg)^{t}$
where $r(x)$ is the number of runs in $x$. Moreover, defining the worst-case insdel ball size as
$B^{\max}_{ID}(k,2t)\triangleq\max_{x\in\mathbb{F}_{2}^{k}}|B_{ID}(x,2t)|,$
we have
$B^{\max}_{ID}(k,2t)\leq\left(\frac{e^{2}(k+t)k}{t^{2}}\right)^{t}.$
Proof.
The first inequality in the upper bound follows from [15, Theorem 4]. For the second inequality, since the number of runs satisfies $r(x)\leq k$. Hence,
$\binom{r(x)+t-1}{t}\sum_{i=0}^{t}\binom{k}{i}\;\leq\;\binom{k+t-1}{t}\sum_{i=0}^{t}\binom{k}{i}$.
Using the standard estimates $\binom{a}{b}\leq\left(\tfrac{ea}{b}\right)^{b}$ and $\sum_{i=0}^{t}\binom{k}{i}\leq\left(\tfrac{ek}{t}\right)^{t}$, we obtain
$\binom{k+t-1}{t}\sum_{i=0}^{t}\binom{k}{i}\;\leq\;\left(\frac{e(k+t)}{t}\right)^{t}\left(\frac{ek}{t}\right)^{t}\;=\;\left(\frac{e^{2}(k+t)k}{t^{2}}\right)^{t}.$
Remark 2.12.
In the literature, an insdel ball is commonly referred to as a fixed-length Levenshtein ball. For a given center word, it is defined as the collection of all sequences of the same length that can be obtained from the center word by performing exactly $t$ deletions and exactly $t$ insertions. Equivalently, this set consists of all words whose insdel distance from the center word is at most $2t$ and whose length is preserved.
Lemma 2.13 (Singleton Bound).
[14] For a binary insdel code $\mathcal{C}$ of length $n$ and minimum distance $d$ , one has
$|\mathcal{C}|\leq 2^{n-d/2+1}$
Lemma 2.14.
Let $u,v\in\mathbb{F}_{2}^{k}$ and $x=uu^{\prime},y=vv^{\prime}\in\mathbb{F}_{2}^{n}$ then $d_{ID}(x,y)\geq d_{ID}(u,v)$.
2.2 Function-Correcting Codes(FCC)
Definition 2.15 ([1]).
[Irregular distance code] Let $\textbf{D}\in\mathbb{N}_{0}^{M\times M}$. A set of binary codewords $\mathcal{P}=\{p_{1},p_{2},\dots,p_{M}\}$ is called an irregular distance code if there exists an ordering of the codewords such that $d_{H}(p_{i},p_{j})\geq[\textbf{D}]_{ij}\quad\text{for all }i,j\in[M].$ We define $N(\textbf{D})$ to be the smallest integer $r$ for which there exists a D-code of length $r$. If $\textbf{D}]_{ij}=D$ for all $i\neq j$, we write $N(M,D)$.
Lemma 2.16 ([1]).
[Plotkin-like bound for irregular distance code] For any distance matrix $\textbf{D}\in\mathbb{N}_{0}^{M\times M}$,
$N(\textbf{D})\geq\begin{cases}\dfrac{4}{M^{2}}\displaystyle\sum_{i,j;\,i<j}[\textbf{D}]_{ij},&\text{if $M$ is even},\\[11.99998pt] \dfrac{4}{M^{2}-1}\displaystyle\sum_{i,j;\,i<j}[\textbf{D}]_{ij},&\text{if $M$ is odd}.\end{cases}$
Lemma 2.17 ([16]).
Let $f:\mathbb{F}_{2}^{k}\rightarrow\mathrm{Im}(f)$ be a locally $(\rho,\lambda)_{\mathrm{H}}$- function. Assume that $\mathrm{Im}(f)$ is equipped with a total order $\prec$, and that for every $u\in\mathbb{F}_{2}^{k}$, the set $B_{\mathrm{H}}^{f}(u,\rho)$ forms a contiguous block with respect to $\prec$. Then there exists a mapping
$\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[\lambda]$
such that for all $u,v\in\mathbb{F}_{2}^{k}$ satisfying $f(u)\neq f(v)\quad\text{and}\quad d_{H}(u,v)\leq\rho$,we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$.
3 Function-Correcting Codes for Insertion-Deletion Errors
In this section, we consider a function defined over the binary vector space $\mathbb{F}_{2}^{k}$, that is, $f:\mathbb{F}_{2}^{k}\rightarrow{Im}(f)$, where the expressiveness of the function is given by E=|Im(f)|E=|{Im}(f)|. Let $x\in\mathbb{F}_{2}^{k}$ denote a binary vector on which the function $f$ is evaluated. The vector $x$ is transmitted over an asynchronous channel that may introduce at most $t$ errors, consisting of insertions, deletions, or both. To enable correction of such errors, $x$ is first encoded using an encoding function $\psi(x)=(x,p(x))$, where $p(x)\in\mathbb{F}_{2}^{r}$ represents the redundancy added to $x$ prior to transmission. Based on this encoding framework, we define function-correcting deletion codes, function-correcting insertion codes, and function-correcting insertion-deletion (insdel) codes, which are designed to recover the function value $f(x)$ in the presence of deletion, insertion, or combined insertion-deletion errors, respectively.
Definition 3.1 (Function-Correcting Deletion Codes(FCDCs)).
An encoding function $\psi(x):\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ with $\psi(x)=(x,p(x))$, $x\in\mathbb{F}_{2}^{k}$ defines a function-correcting deletion codes for $f:\mathbb{F}_{2}^{k}\rightarrow{Im}(f)$ if for all $x$ and $y$ such that $f(x)\neq f(y)$, following holds
$D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset$
The aforementioned formulation guarantees that, even after up to $t$ deletions, the associated codewords for any two inputs $x$ and $y$ that result in different function values $f(x)$ and $f(y)$ remain recognizable. Consequently, any received subsequence resulting from at most $t$ deletions can be uniquely mapped back to a codeword, and hence to its corresponding function value $f(x)$. On the same line, function-correcting insertion codes can be defined as follows:
Definition 3.2 (Function-Correcting Insertion Codes(FCICs)).
An encoding function $\psi(x):\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ with $\psi(x)=(x,p(x))$, $x\in\mathbb{F}_{2}^{k}$ defines a function-correcting insertion codes for $f:\mathbb{F}_{2}^{k}\rightarrow\operatorname{Im}(f)$ if for all $x$ and $y$ such that $f(x)\neq f(y)$, following holds
$I_{t}(\psi(x))\cap I_{t}(\psi(y))=\emptyset$
Function-correcting insdel codes, designed to recover function evaluations in the presence of both insertion and deletion errors, are defined as follows.
Definition 3.3 (Function-Correcting Insdel Codes(FCIDCs)).
The encoding map $\psi:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$ defined as $\psi(x)=(x,p(x))$ yields a function-correcting code for the function $f:\mathbb{F}_{2}^{k}\rightarrow\text{Im}(f)$ in insdel metric if for all $x$ and $y$ such that $f(x)\neq f(y)$, the insdel distance $d_{ID}(\psi(x),\psi(y))>2t$.
This condition ensures that up to $t$ insertions and deletions can be corrected to recover the function value $f(x)$ from any corrupted version of $\psi(x)$. The next proposition establishes equivalence between the above-defined codes.
Proposition 3.4.
Let $\psi:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{n}$ be an encoding map with $n=k+r$, and fix $t\geq 1$. The following conditions are equivalent for distinct inputs $x,y\in\mathbb{F}_{2}^{k}$ with $f(x)\neq f(y)$:
  1. 1.
    $D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset$.
  2. 2.
    $I_{t}(\psi(x))\cap I_{t}(\psi(y))=\emptyset$.
  3. 3.
    $d_{ID}(\psi(x),\psi(y))>2t$.
In particular, a code that is function-correcting for up to $t$ deletions(insertions) is also function-correcting for up to $t$ insertion-deletions.
Proof.
Let $n=|\psi(x)|=|\psi(y)|$.
(1) $\Rightarrow$ (3). If $D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset$, then there is no common subsequence of length at least ntn-t for $\psi(x)$ and $\psi(y)$. Hence
$\operatorname{LCS}(\psi(x),\psi(y))<n-t,$
So $d_{ID}(\psi(x),\psi(y))=2(n-\operatorname{LCS}(\psi(x),\psi(y)))>2t$. Hence every $t$-function-correcting deletion code is $t$-function-correcting insdel code.
(3) $\Rightarrow$ (1). If $d_{ID}(\psi(x),\psi(y))>2t$, then $\operatorname{LCS}(\psi(x),\psi(y))<n-t$. Hence, there cannot exist a common subsequence of length $\geq n-t$. Therefore, we get
$(D_{t}(\psi(x))\cap D_{t}(\psi(y))=\emptyset.$
(1) $\Rightarrow$ (2).(By contradiction) Suppose that there exists a word $w$ with $w\in I_{t}(\psi(x))\cap I_{t}(\psi(y))$. Then both $\psi(x)$ and $\psi(y)$ are subsequences of $w$. Let $|w|=n+s$ with $s\leq t$. Embedding two length-$n$ sequences into a common supersequence of length $n+s$ forces their LCS to be at least nsn-s. Therefore,
$\operatorname{LCS}(\psi(x),\psi(y))\geq n-s\geq n-t,$
This implies the existence of a common subsequence of length $\geq n-t$. This contradicts (1). Thus $I_{t}(\psi(x))\cap I_{t}(\psi(y))=\emptyset$.
[…]
Given the equivalence among the three code formulations discussed above, we restrict our attention, in the rest of this paper, to function-correcting insertion-deletion codes (FCIDCs) .
Next, we define the optimal redundancy of an FCIDC for a function $f$ as it is a key parameter and plays a pivotal role in the study of FCIDCs.
Definition 3.5 (optimal redundancy).
A positive integer $r$ is called the optimal redundancy of a function-correcting insdel code for a function $f$, defined by an encoding map
$\psi:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r},$
if $r$ is the smallest achievable redundancy length. The optimal redundancy is denoted by $r^{f}_{ID}(k,t)$.
3.1 Irregular Insdel-Distance Codes
In this section, we define irregular insdel-distance codes and establish their relationship with FCIDCs. Leveraging this connection, we derive several general results about FCIDCs and obtain both lower and upper bounds on their optimal redundancy.
We first define insdel distance matrices (Definition 3.6) for a function $f$, followed by irregular insdel-distance codes (Definition 3.10) in which the insdel-distance between each pair of codewords should satisfy individual distance constraints.
Definition 3.6 (Insdel Distance Matrices).
Let $M,t\in\mathbb{N}$. Consider $M$ binary vectors ${x_{1}},...,{x_{M}}\in{\mathbb{F}^{k}_{2}}$. Then, ${\textbf{I}_{f}^{(1)}}(t,{x_{1}},...,{x_{M}})$ and ${\textbf{I}_{f}^{(2)}}(t,x_{1},\ldots,x_{M})$ are $M\times M$ insdel distance matrices corresponding to function $f$ with entries as follows:
[…]
and
[…]
Remark 3.7.
Unlike the Hamming metric, the insdel metric allows cross-block alignment (Lemma 3.14) between concatenated words, leading to a loss of up to $2\min\{k,r\}$ in insdel distance. The additional $2k$ term in distance matrix $\bm{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})]$ precisely compensates for this phenomena when $r\geq k.$
The following example illustrates that the +2k+2k term in the definition for $\bm{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})$ is indispensible.
Example 3.8.
Let $k=4$ and $t=4$, and let $f:\mathbb{F}_{2}^{4}\to\mathrm{Im}(f)$ be a function such that $f(0000)\neq f(1111)$. Since
$d_{ID}(0000,1111)=2(4-0)=8,$
the distance requirement on the redundancy vectors corresponding to matrix $\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})$ reduces to
$d_{ID}\bigl(p(x_{1}),p(x_{2})\bigr)\;\geq\;2t+2-d_{ID}(x_{1},x_{2})=2.$
Choose the redundancy vectors
$p(x_{1})=101110,\qquad p(x_{2})=000000,$
for which
$d_{ID}\bigl(p(x_{1}),p(x_{2})\bigr)=8,$
and hence the $\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})$ condition is satisfied. Consider the concatenated codewords
$\psi(x_{1})=0000101110,\qquad\psi(x_{2})=1111000000.$
The longest common subsequence between these two words has length $6$, yielding
$d_{ID}\bigl(\psi(x_{1}),\psi(x_{2})\bigr)=2(10-6)=8<2t+2.$
Thus, despite satisfying the $\mathbf{I}_{f}^{(1)}(t,x_{1},\dots,x_{M})$ distance requirement, the resulting code fails to correct $t=4$ insertion–deletion errors.
In contrast, the distance requirement for the matrix $\mathbf{I}_{f}^{(2)}(t,x_{1},\dots,x_{M})$ enforces
$d_{ID}\bigl(p(x_{1}),p(x_{2})\bigr)\;\geq\;2t+2+2k-d_{ID}(x_{1},x_{2})=10,$
which excludes this choice of redundancy vectors. This example demonstrates that the additional $2k$ compensation term is essential to prevent distance collapse caused by cross LCS effects.
An illustrative example corresponding to each of the two types of matrices is presented next.
Example 3.9.
Let $f:\{0,1\}^{2}\to\{0,1\}$ be given by
$f(10)=f(01)=0\text{ and }f(00)=f(11)=1.$
Fix the input ordering
$(x_{1},x_{2},x_{3},x_{4})=(00,01,10,11).$
For $t=1$, the matrices $\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4})$ and $\textbf{I}_{f}^{2}(t;x_{1},x_{2},x_{3},x_{4})$ are
$\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4})=\begin{pmatrix}0&2&2&0\\ 2&0&0&2\\ 2&0&0&2\\ 0&2&2&0\end{pmatrix}$,
$\textbf{I}_{f}^{2}(t;x_{1},x_{2},x_{3},x_{4})=\begin{pmatrix}0&6&6&0\\ 6&0&0&6\\ 6&0&0&6\\ 0&6&6&0\end{pmatrix}$.
Let $\mathcal{P}=\{p_{1},p_{2},\dots,p_{M}\}\subseteq\mathbb{F}_{2}^{r}$ be a code of size $M$ and length $r$. The unconventional choice of using the code-block length $r$ is motivated by its relationship to the redundancy of FCIDCs, which is discussed later in this section.
Definition 3.10 (Irregular insdel-distance codes).
Let I be a square matrix of order $M$ whose entries are non-negative integers, and let $K\in\mathbb{N}$. Then,
  • $\mathcal{P}$ is an irregular insdel-distance code of type $1$ for matrix $\mathbf{I}$ if there exists an ordering of $\mathcal{P}$ such that
    $d_{ID}(p_{i},p_{j})\geq[\bm{I}]_{i,j}\quad\text{for all }1\leq i,j\leq M.$
  • $\mathcal{P}$ is an irregular insdel-distance code of type $2$ corresponding to the matrix $\mathbf{I}$ if it is an irregular insdel-distance code of type $1$ and, in addition, its codeword length satisfies $r\geq K$.
Next, we define the shortest achievable length for both types of irregular insdel-distance codes.
Definition 3.11.
[…]
When $\bm{I}]_{i,j}=I$ for all $i\neq j$, where $I\in\mathbb{N}$, we write $N^{(1)}_{ID}(M,I)$ and $N^{(2)}_{{ID}}(M,I;K)$, respectively.
Example 3.12.
Let $p_{1}=(0),p_{2}=(1),p_{3}=(1)$ and $p_{4}=(0)$. Then, $\{p_{1},p_{2},p_{3},p_{4}\}$ is an irregular insdel-distance code of type $1$ corresponding to matrix $\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4})$ of Example 3.9. Clearly, $N^{(1)}_{{ID}}(\textbf{I}_{f}^{1}(t;x_{1},x_{2},x_{3},x_{4}))=1.$
Now, let $K=2$ then $p_{1}=(000),p_{2}=(111),p_{3}=(111)$ and $p_{4}=(000)$ is an irregular insdel-distance code of type $2$ corresponding to matrix $\textbf{I}_{f}^{(2)}(t;x_{1},x_{2},x_{3},x_{4})$ of Example 3.9. In this case $N^{(2)}_{{ID}}(\textbf{I}_{f}^{(2)}(t;x_{1},x_{2},x_{3},x_{4});2)=3.$
The results in the next two lemmas play a crucial role in establishing the relationship between the optimal redundancy of FCIDCs and the lengths of irregular insdel-distance codes.
Lemma 3.13.
Let $x=(x_{1},x_{2})\in\mathbb{F}_{2}^{k+r}$ and $y=(y_{1},y_{2})\in\mathbb{F}_{2}^{k+r}$ where $x_{1},y_{1}\in\mathbb{F}_{2}^{k}$ and $x_{2},y_{2}\in\mathbb{F}_{2}^{r}$ then,
[…]
Proof.
Let $s_{1}$ be the longest common subsequence of $x_{1}$ and $y_{1}$, and $s_{2}$ be the longest common subsequence of $x_{2}$ and $y_{2}$; then $s_{1}s_{2}$ is a common subsequence of $x$ and $y$. Therefore,
$LCS(x,y)\geq LCS(x_{1},y_{1})+LCS(x_{2},y_{2})$
For the right-hand inequality, let $s$ be the longest common subsequence of $x$ and $y$. For each symbol in $s$ consider the positions to which it is matched in both $x$ and $y$. This induces a partition of the symbols of $s$ into four classes indexed by $(i,j),\quad i,j\in\{1,2\}$, where a symbol is of type $(i,j)$ if it is matched to $x_{i}$ in $x$ and to $y_{j}$ in $y$. Let $n_{ij}$ denote the number of symbols of type $(i,j)$ in $s$. Then following four cases can be considered:
Case 1: If the symbols are of type $(1,1)$ then they form a common subsequence of $x_{1}$ and $y_{1}$.Therefore,
$n_{11}\leq LCS(x_{1},y_{1}).$
Case 2: If the symbols are of type $(2,2)$ then they form a common subsequence of $x_{2}$ and $y_{2}$.Therefore,
$n_{22}\leq LCS(x_{2},y_{2}).$
Case 3: If the symbols are of type $(1,2)$ then they form a common subsequence of $x_{1}$ and $y_{2}$.Therefore,
$n_{12}\leq\min\{k,r\}$
Case 4: If the symbols are of type $(2,1)$ then they form a common subsequence of $x_{2}$ and $y_{1}$.Therefore,
$n_{21}\leq\min\{k,r\}$
We claim $n_{12}\cdot n_{21}=0,$ i.e., symbols of type $(1,2)$ and $(2,1)$ cannot appear together in $s$. On the contrary assume that there exists symbols $s_{i}$ and $s_{j}$ of type $(1,2)$ and $(2,1)$ of $s$, if $s_{i}$ preceeds $s_{j}$, then the ordering is respected in $x$ but violated in $y$. Similarly, if $s_{j}$ preceeds $s_{i}$, then the ordering is respected in $y$ but violated in $x$.
Hence, the claim follows and we have $n_{12}+n_{21}=\max\{n_{12},n_{21}\}\leq\min\{k,r\}.$
Lemma 3.14.
Let $x=(x_{1},x_{2})\in\mathbb{F}_{2}^{k+r}$ and $y=(y_{1},y_{2})\in\mathbb{F}_{2}^{k+r}$ where $x_{1},y_{1}\in\mathbb{F}_{2}^{k}$ and $x_{2},y_{2}\in\mathbb{F}_{2}^{r}$ then,
[…]
Proof.
From the definition of the insdel-distance, we have
$d_{ID}(x,y)=2(k+r)-2LCS(x,y)$
$d_{ID}(x_{1},y_{1})=2k-LCS(x_{1},y_{1})$
$d_{ID}(x_{2},y_{2})=2r-LCS(x_{2},y_{2})$
multiplying (2)(-2) to equation 3.1 we get:
$-2(LCS(x_{1},y_{1})+LCS(x_{2},y_{2})+\min\{k,r\}$ $\leq-2\cdot LCS(x,y)$
$\leq-2(LCS(x_{1},y_{1})+LCS(x_{2},y_{2}))$
Adding $2(k+r)$ to each part in the above inequality, we get
$2(k+r)-2(LCS(x_{1},y_{1})+LCS(x_{2},y_{2})+\min\{k,r\})$ $\leq 2(k+r)-2\cdot LCS(x,y)$
$\leq 2(k+r)-2(LCS(x_{1},y_{1})+LCS(x_{2},y_{2}))$
$2k-2\cdot LCS(x_{1},y_{1})+2r-LCS(x_{2},y_{2})-2\cdot\min\{k,r\}$ $\leq 2(k+r)-2LCS(x,y)$
$\leq 2k-2\cdot LCS(x_{1},y_{1})+2r-LCS(x_{2},y_{2})$
Therefore,
$d_{ID}(x_{1},y_{1})+d_{ID}(x_{2},y_{2})-2\cdot\min\{k,r\}\leq d_{ID}(x,y)\leq d_{ID}(x_{1},y_{1})+d_{ID}(x_{2},y_{2})$
Remark 3.15.
Lemma 3.14 suggests that cross-LCS effects cause the insdel distance between concatenated codewords to suffer a worst-case loss of $2\min\{k,r\}$ for redundancy length $r$. The resulting distance condition simplifies to demanding that the message vectors themselves already satisfy the full insdel distance constraint if $r<k$. This loss completely offsets any redundancy distance gain. Determining the type $2$ irregular insdel-distance code is motivated by the fact that nontrivial systematic function-correcting insertion-deletion codes only exist in the regime $r\geq k$.
The following example shows that the lower bound given in Lemma 3.14 is tight.
Example 3.16.
Let $x=(0000,10111)\in\mathbb{F}_{2}^{9}$ and $y=(1111,00000)\in\mathbb{F}_{2}^{9}$. Then,
$d_{ID}(0000,1111)=8,\quad d_{ID}(10111,00000)=8\text{ and }d_{ID}(000010111,111100000)=8$.
Hence, $d_{ID}(000010111,111100000)=d_{ID}(0000,1111)+d_{ID}(10111,00000)-2\cdot 4$.
The following theorem establishes upper and lower bounds on the optimal redundancy of FCIDCs constructed for generic functions. Moreover, it reduces the computation of $r^{f}_{ID}(k,t)$ to determining the quantities $N_{ID}^{(1)}\left(\bm{I}^{(1)}_{f}(t,x_{1},\ldots,x_{2^{k}})\right)$ and $N_{ID}^{(2)}\!\left(\bm{I}^{(2)}_{f}(t,u_{1},\ldots,u_{2^{k}})\right)$.
Theorem 3.17.
For any function $f:\mathbb{F}^{k}_{2}\rightarrow Im(f)$ we have
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))\leq r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k).$
Proof.
We establish the theorem by considering the following cases.
Case 1: (Constant functions) If $f$ is a constant function, then $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},\ldots,x_{2^{k}}))=0$ and $N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},\ldots,x_{2^{k}};k))=k$. Consequently, the desired condition holds trivially, establishing the theorem in this case. Specifically, we have:
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t))=r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t)).$
Case 2: (Non-constant functions) To prove that $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))\leq r_{ID}^{f}(k,t)$ for a non-constant function $f$, consider a function-correcting insdel code for $f$ defined by an encoding function
$\psi:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{k+r},\quad\bm{x_{i}}\mapsto(\bm{x_{i}},{p_{i}}),$
where the redundancy $r$ is optimal, i.e., $r=r_{ID}^{f}(k,t)$. On the contrary, suppose that
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))>r_{ID}^{f}(k,t)$. This implies the existence of distinct indices $i,j\in\{1,\ldots,2^{k}\}$ such that $f({x_{i}})\neq f({x_{j}})$ and
$d_{ID}({p_{i}},{p_{j}})<2t+2-d_{ID}({x_{i}},{x_{j}}).$
Consequently, from equation 3.2 we obtain
$d_{ID}(\psi({x_{i}}),\psi({x_{j}}))\leq d_{ID}({x_{i}},{x_{j}})+d_{ID}({p_{i}},{p_{j}})<2t+2.$
This contradicts that $\psi$ defines a function-correcting insdel code. Hence,
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))\leq r_{ID}^{f}(k,t).$
Next, we establish the reverse inequality $r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$. Let $\mathcal{P}=\{{p_{1}},\ldots,{p_{2^{k}}}\}$ be a $\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}})$ irregular insdel-distance code of type 2 and length $N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$, and define the encoding function
$\psi:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{k+r},\quad{x_{i}}\mapsto({x_{i}},{p_{i}}).$
For every pair $i,j\in\{1,\ldots,2^{k}\}$ with $f({x_{i}})\neq f({x_{j}})$, we have
$d_{ID}(\psi({x_{i}}),\psi({x_{j}}))$ $\geq d_{ID}({x_{i}},{x_{j}})+d_{ID}({p_{i}},{p_{j}})$
$\quad-2\cdot\min\{k,N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},...,x_{2^{k}};k))\}.$
$\geq d_{ID}({x_{i}},{x_{j}})+d_{ID}({p_{i}},{p_{j}})-2k$
$\geq d_{ID}({x_{i}},{x_{j}})+2t+2+2k-d_{ID}({x_{i}},{x_{j}})$
$\qquad-2k$
$=2t+2.$
Thus, $\psi$ defines a function-correcting insdel code for $f$ with redundancy $r=N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k)$, yielding
$r^{f}_{ID}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,x_{1},x_{2},...,x_{2^{k}});k).$
This completes the proof. ∎
Example 3.18.
With the help of Example 3.12 and Theorem 3.17 we can bound the optimal redundancy of function $f$ defined in Example 3.9 as follows:
$1\leq r_{ID}^{f}(2,1)\leq 3.$
3.2 Simplified Redundancy Lower Bounds
Using a smaller set of information vectors, one can obtain a lower bound on the redundancy of FCIDCs as follows:
Corollary 3.19.
Consider a collection of $M$ distinct binary words $x_{1},x_{2},\ldots,x_{M}$ of length $k$. Then, for a function $f$, the optimal redundancy of an FCIDCs is lower bounded as follows:
$r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})).$
For a non-constant function $f$,
$r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(\bm{I}_{f}^{(2)}(2,2t-1))=t.$
Proof.
Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{2^{k}}\}$ be $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}})$-code of length $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}}))$ then $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$ is a $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})$-code. Hence,
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M}))\leq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{2^{k}})).$
From Theorem 3.17 we get $r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},...,x_{M})).$
Since $E\geq 2$ for a non-constant function and from [1, Corollary 1] we know that there exists $x,x^{\prime}\in\mathbb{F}_{2}^{k}$ with $d_{H}(x,x^{\prime})=1$ and $f(x)\neq f(x^{\prime})$ therefore using the inequality, $2\leq d_{ID}(x,x^{\prime})\leq 2d_{H}(x,x^{\prime})$ we get $d_{ID}(x,x^{\prime})=2$, that is there will always exist $x,x^{\prime}\in\mathbb{F}_{2}^{k}$ with $d_{ID}(x,x^{\prime})=2$ such that $f(x)\neq f(x^{\prime})$ whenever $E\geq 2.$ So, in particular for $M=2,$ we can say $r^{f}_{ID}(k,t)\geq N^{(1)}_{ID}(2,2t).$ Consider the following repetition code of length $t$, $\mathcal{C}=\{(0,0,\ldots,0),(1,1,\ldots,1)\}$ then $d_{ID}(\mathcal{C})=2t$. Hence, $N^{(1)}_{ID}(2,2t)=t.$
As pointed out in [1], finding the optimal length of irregular distance codes over a complete set of message vectors is quite difficult. In order to obtain a more computationally easier bound as compared to the Theorem 3.17 we define the concept of function distance and function distance matrices.
Definition 3.20 (Function Distance).
Let $f_{1},f_{2}\in Im(f)$, then the minimum insdel distance between two information vectors that evaluate to $f_{1}$ and $f_{2}$ gives the insdel distance between the function values $f_{1}$ and $f_{2}$, i.e.,
$d_{ID}^{f}(f_{1},f_{2})=\min_{x_{1},x_{2}\in\mathbb{F}_{2}^{k}}d_{ID}(x_{1},x_{2})\quad\text{s.t.}\text{ }f(x_{1})=f_{1}\text{ and }f(x_{2})=f_{2}.$
Definition 3.21 (Function Distance Matrices).
The function insdel-distance matrices for function $f$ are square matrices of order $E=|\text{Im}(f)|$ denoted by $\bm{I}_{f}^{(1)}(t,f_{1},\ldots,f_{E})$ and $\bm{I}_{f}^{(2)}(t,f_{1},\ldots,f_{E})$ respectively and whose entries are given by:
$[\bm{I_{f}^{(1)}}(t,f_{1},\dots,f_{E})]_{ij}=\begin{cases}[2(t+1)-d_{ID}^{f}(f_{i},f_{j})]^{+},&\text{if }i\neq j,\\ 0,&\text{otherwise}.\end{cases}$
and
$[\bm{I_{f}^{(2)}}(t,f_{1},\dots,f_{E})]_{ij}=\begin{cases}[2(t+1+k)-d_{ID}^{f}(f_{i},f_{j})]^{+},&\text{if }i\neq j,\\ 0,&\text{otherwise}.\end{cases}$
Theorem 3.22.
For arbitrary function $f:\mathbb{F}_{2}^{k}\to\mathrm{Im}(f)$, we have
$r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}\big(\bm{I}_{f}^{(2)}(t,f_{1},\ldots,f_{E});k\big).$
Proof.
Consider the following construction method for FCIDCs. Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{E}\}$ be a set of codewords of length $r$ such that $r\geq k$ and $d_{ID}(p_{i},p_{j})\geq 2t+2+2k-d_{ID}^{f}(f_{i},f_{j}),\quad\forall i,j\in[E]$. Let $p_{i}$ be the redundancy vector corresponding to all $u\in\mathbb{F}_{2}^{k}$ such that $f(u)=f_{i}$, i.e., all message vectors mapping to the same function value have the same redundancy vector. From 3.14 it follows that for any $\mathbf{u}_{i},\mathbf{u}_{j}\in\mathbb{F}_{2}^{k}$ with $f(\mathbf{u}_{i})=f_{i}$, $f(\mathbf{u}_{j})=f_{j}$, $f_{i}\neq f_{j}$, we have
$d_{ID}(\psi(\mathbf{u}_{i}),\psi(\mathbf{u}_{j}))$ $\geq d_{ID}(\mathbf{u}_{i},\mathbf{u}_{j})+d_{ID}(p_{i},p_{j})-2\cdot min\{k,r\}$
$\geq d_{ID}(u_{i},u_{j})+2t+2+2k-d_{ID}^{f}(f_{i},f_{j})-2k$
$\geq 2t+2.$
According to the definition of FCIDCs, we have constructed an FCIDC. Hence,
$r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}\big({\bm{I}}_{f}^{(2)}(t,f_{1},\ldots,f_{E});k\big).$
It is not always easy to derive a generic formula for the function distance $d_{ID}^{f}(f_{i},f_{j})$. Consequently, an upper bound on the optimal redundancy can be estimated using a lower constraint on $d_{ID}^{f}(f_{i},f_{j})$ in the absence of an explicit equation.
Lemma 3.23.
Let $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a function with expressivness $E=|Im(f)|$ and let $\big\{a_{ij}\big\}_{1\leq i,j\leq E}$ be a set of $E$ non-negative integers such that $a_{ij}\leq d_{ID}^{f}(f_{i},f_{j})$ for all $i,j$. Then,
$N_{ID}^{(2)}(\bm{I}_{f}^{(2)}(t,f_{1},f_{2},...,f_{E});k)\leq N_{ID}^{(2)}(I;k)$
where $I$ is a symmetric square matrix of order $E$ whose entries are given by
$[{I}]_{ij}=\begin{cases}[2t+2+2k-a_{ij}]^{+},&\text{if }i\neq j,\\ 0,&\text{otherwise}.\end{cases}$
Proof.
Let $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{E}\}$ be a length $N_{ID}^{(2)}(I;k)$, irregular insdel-distance code of type $2$ corresponding to matrix $I$. Then, for all $i\neq j$
$d_{ID}(c_{i},c_{j})$ $\geq[2t+2+2k-a_{ij}]^{+}$
$\geq[2t+2+2k-d_{ID}(f_{i},f_{j})]^{+}$
$=[\bm{I}_{f}^{(2)}(t,f_{1},\dots,f_{E})]_{ij}$
Hence, $\mathcal{C}$ is an irregular insdel-distance code of type $2$ for function distance matrix ${\bm{I}_{f}^{(2)}}(t,f_{1},\ldots,f_{E})$. Therefore, $N_{ID}^{(2)}(\bm{I}_{f}^{(2)}(t,f_{1},f_{2},...,f_{E});k)\leq N_{ID}^{(2)}(I;k).$
Combining Lemma 3.23 and Theorem 3.22 yields the following corollary.
Corollary 3.24.
Let $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a function with expressivness $E=|Im(f)|$ and let $\big\{a_{ij}\big\}_{1\leq i,j\leq E}$ be a set of $E$ non-negative integers such that $a_{ij}\leq d_{f}(f_{i},f_{j})$ for all $i,j$. Then,
$r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(I;k)$
where $I$ is a symmetric square matrix of order $E$ whose entries are given by
$[{I}]_{ij}=\begin{cases}[2t+2+2k-a_{ij}]^{+},&\text{if }i\neq j,\\ 0,&\text{otherwise}.\end{cases}$
The next lemma is derived from the proof of [17, Lemma 1] by assuming the existence of a common binary super-sequence $v$ of length $N$ for all codewords in the binary insdel code, and setting $t_{D}=0$ (number of deletions) and $t_{I}=N-n$ (number of insertions), where $n$ is the length of the codewords.
Lemma 3.25.
[17] Let $C\subseteq\mathbb{F}_{2}^{r}$ be a code of minimum insdel distance $d$ and size $M$. Suppose that there exists $v\in\mathbb{F}_{2}^{N}$ that contains every codeword in $C$ as a subsequence, then
$\sum_{c_{i},c_{j}\in C}d_{ID}(c_{i},c_{j})\leq\frac{M^{2}(N-n)n}{N}$
Using the result from the previous lemma, we will now establish a lower bound on $N_{ID}^{(1)}(I)$ for a given distance matrix $I$. This bound can be viewed as a generalization of the Plotkin-type bound for insdel codes with irregular distance requirements.
Lemma 3.26.
For $N\in\mathbb{N}$ and any matrix $\bm{I}\in\mathbb{N}_{0}^{M\times M}$, such that $N\geq\frac{4S}{M^{2}}$,
$N_{ID}^{(1)}(I)\geq\frac{N-\sqrt{N^{2}-\frac{4SN}{M^{2}}}}{2}$
where $S=\sum_{i,j:i<j}[I]_{ij}$
Proof.
Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$ be a length $r$, type $1$ irregular insdel-distance code corresponding to matrix $I$ and let $v$ be an $N$-length binary sequence such that every codeword in $\mathcal{P}$ is a subsequence of $v$. Then from Lemma 3.25, $\sum_{i,j;i<j}d_{ID}(p_{i},p_{j})\leq\frac{M^{2}(N-r)r}{N}$. Since,
$S=\sum_{i,j:i<j}[I]_{ij}\leq\sum_{i,j;i<j}d_{ID}(p_{i},p_{j})\leq\frac{M^{2}(N-r)r}{N}$
we have
$N_{ID}^{(1)}(I)=r\geq\frac{N-\sqrt{N^{2}-\frac{4SN}{M^{2}}}}{2}.$
Corollary 3.27.
For $N\in\mathbb{N}$ and any matrix $\bm{I}\in\mathbb{N}_{0}^{M\times M}$, such that $N=\frac{4\sum_{i,j:i<j}[I]_{ij}}{M^{2}}$,
$N_{ID}^{(1)}(I)\geq\frac{2}{M^{2}}\sum_{i,j:i<j}[I]_{ij}$
The result in Corollary 3.27 can also be derived from the plotkin-like bound for codes with irregular distance requirement given in [1] for the Hamming metric, together with the observation that if the insdel distance between two codewords is $d$, then their Hamming distance is at least $\frac{d}{2}$. For completeness, we present below a lemma that formally proves this result using the aforementioned approach.
Lemma 3.28.
For any insdel matrix $\bm{I}\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{ID}^{(1)}(\bm{I)}\;\geq\;\begin{cases}\dfrac{2}{M^{2}}\displaystyle\sum_{i<j}[I]_{ij},&\text{if $M$ is even},\\[11.99998pt] \dfrac{2}{M^{2}-1}\displaystyle\sum_{i<j}[I]_{ij},&\text{if $M$ is odd}.\end{cases}$
Proof.
For any $i<j$, we have from Lemma 2.8
$d_{H}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\,d_{ID}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\,[I]_{ij},$
summing over all pairs yields
$\sum_{i<j}d_{H}(p_{i},p_{j})\;\geq\;\tfrac{1}{2}\sum_{i<j}[I]_{ij}.$
From Lemma 2.16, we know in case of even $M$
$\sum_{i<j}d_{H}(p_{i},p_{j})\;\leq\;r\cdot\big\lfloor M^{2}/4\big\rfloor.$
and when $M$ is odd,
$\sum_{i<j}d_{H}(p_{i},p_{j})\;\leq\;r\cdot\big\lfloor(M-1)^{2}/4\big\rfloor.$
Combining $(1)$ and $(2)$ and then $(1)$ and $(3)$ gives
$r\cdot\big\lfloor M^{2}/4\big\rfloor\;\geq\;\tfrac{1}{2}\sum_{i<j}[I]_{ij},$
$r\cdot\big\lfloor(M-1)^{2}/4\big\rfloor\;\geq\;\tfrac{1}{2}\sum_{i<j}[I]_{ij},$
and rearranging yields the stated bound ∎
The next lemma is a variant of the Gilbert-Varshamov bound for irregular insdel-distance codes.
Lemma 3.29.
For any distance matrix $\bm{I}\in\mathbb{N}^{M\times M}_{0}$ and any permutation $\pi:\{1,2,\ldots,M\}\rightarrow\{1,2,\ldots,M\}$, we have
$N^{(1)}_{ID}(\bm{I})\leq\min_{r\in\mathbb{N}}\left\{r\ \middle|\ 2^{r}>\max_{j\in\{1,2,\ldots,M\}}\sum_{i=1}^{j-1}B^{\max}_{ID}\big(r,[\bm{I}]_{\pi(i)\pi(j)}-2\big)\right\},$
and
$N^{(2)}_{ID}(\bm{I};K)\leq\min_{r\geq K}\left\{r\ \middle|\ 2^{r}>\max_{j\in\{1,2,\ldots,M\}}\sum_{i=1}^{j-1}B^{\max}_{ID}\big(r,[\bm{I}]_{\pi(i)\pi(j)}-2\big)\right\}.$
$B^{\max}_{ID}(r,t)\triangleq\max_{x\in\mathbb{F}_{2}^{r}}\left|B_{ID}(x,t)\right|,\qquad B_{ID}(x,t)=\{\,y\in\mathbb{F}_{2}^{r}\mid d_{ID}(x,y)\leq t\,\}.$
Proof.
We present an iterative way of selecting valid codewords for constructing an irregular insdel-distance code of type 2 (type 1) for any distance matrix $\bm{I}$. For simplicity, first choose $\pi$ as the identity permutation. Select an arbitrary vector $p_{1}\in\mathbb{F}_{2}^{r}$ as the first codeword, where $r\geq K$ (or $r\in\mathbb{N}$ for type 1). For the second codeword $p_{2}$, the distance requirement
$d_{ID}(p_{1},p_{2})\geq[\bm{I}]_{12}$
must be satisfied. Such a $p_{2}$ exists provided
$2^{r}>\left|B_{ID}\big(p_{1},[\bm{I}]_{12}-2\big)\right|.$
Since the size of an insdel ball depends on its center, we upper bound it by the worst-case ball size $B^{\max}_{ID}(r,[\bm{I}]_{12}-2)$.
Next, choose the third codeword $p_{3}$ such that
$d_{ID}(p_{1},p_{3})\geq[\bm{I}]_{13}\quad\text{and}\quad d_{ID}(p_{2},p_{3})\geq[\bm{I}]_{23}.$
A sufficient condition for the existence of such a $p_{3}$ is
$2^{r}>\left|B_{ID}\big(p_{1},[\bm{I}]_{13}-2\big)\right|+\left|B_{ID}\big(p_{2},[\bm{I}]_{23}-2\big)\right|$
which is guaranteed if
$2^{r}>B^{\max}_{ID}\big(r,[\bm{I}]_{13}-2\big)+B^{\max}_{ID}\big(r,[\bm{I}]_{23}-2\big).$
Proceeding inductively, at the $j$-th step we must choose $p_{j}$ such that
$d_{ID}(p_{i},p_{j})\geq[\bm{I}]_{ij}\quad\text{for all }i<j.$
A sufficient condition for the existence of $p_{j}$ is
$2^{r}>\sum_{i=1}^{j-1}B^{\max}_{ID}\big(r,[\bm{I}]_{ij}-2\big).$
Since the codewords can be selected in any order, the same argument holds for any permutation $\pi$, completing the proof. ∎
Theorem 3.30.
For a positive integer $M$, an even positive integer $d$, and $K\geq 2$, we have
$N_{ID}^{(2)}(M,d;K)\leq\Bigg\lceil\frac{\ln M+(d-2)\ln\!\Big(\frac{2eK}{d-2}\Big)}{\ln 2}\Bigg\rceil.$
Proof.
By Lemma 3.29, the minimum length of a type 2 irregular insdel-distance code with $M$ codewords and distance requirement $d$ satisfies
$N^{(2)}_{ID}(M,d;K)$ $\leq\min_{r\geq K}\left\{r\,\middle|\,2^{r}>\max_{j\in\{1,\ldots,M\}}\sum_{i=1}^{j-1}B^{\max}_{ID}(r,d-2)\right\}$
$\leq\min_{r\geq K}\left\{r\,\middle|\,2^{r}>M\,B^{\max}_{ID}(r,d-2)\right\},$
where
$B^{\max}_{ID}(r,t)\triangleq\max_{x\in\mathbb{F}_{2}^{r}}\left|B_{ID}(x,t)\right|.$
Using the upper bound on the insdel ball size from Lemma 2.11, we have
$B^{\max}_{ID}(r,d-2)\leq\frac{e^{d-2}\,r^{\frac{d-2}{2}}\,\left(r+\frac{d-2}{2}\right)^{\frac{d-2}{2}}}{\left(\frac{d-2}{2}\right)^{d-2}}.$
Let $d^{\prime}=d-2$. A sufficient condition for $M\cdot\,B^{\max}_{ID}(r,d^{\prime})<2^{r}$ is
$\frac{e^{d^{\prime}}\,r^{\frac{d^{\prime}}{2}}\,\left(r+\frac{d^{\prime}}{2}\right)^{\frac{d^{\prime}}{2}}}{\left(\frac{d^{\prime}}{2}\right)^{d^{\prime}}}\,M<2^{r}.$
Taking natural logarithms on both sides yields
$d^{\prime}+\frac{d^{\prime}}{2}\ln r+\frac{d^{\prime}}{2}\ln\!\left(r+\frac{d^{\prime}}{2}\right)-d^{\prime}\ln\!\left(\frac{d^{\prime}}{2}\right)+\ln M<r\ln 2.$
Using $d^{\prime}\ln\!\left(\frac{d^{\prime}}{2}\right)=-d^{\prime}\ln d^{\prime}+d^{\prime}\ln 2$ and the inequality $\ln\!\left(r+\frac{d^{\prime}}{2}\right)\geq\ln r,$ we obtain
$\ln M+d^{\prime}-d^{\prime}\ln d^{\prime}+d^{\prime}\ln 2<r\ln 2-d^{\prime}\ln r.$
Since $r\geq K$, we have $\ln r\geq\ln K$, and hence
$r\ln 2>\ln M+d^{\prime}+d^{\prime}\ln\!\Big(\frac{2K}{d^{\prime}}\Big).$
Dividing both sides by $\ln 2$ gives
$r>\frac{\ln M+d^{\prime}\ln\!\Big(\frac{2eK}{d^{\prime}}\Big)}{\ln 2}.$
Therefore,
$N_{ID}^{(2)}(M,d;K)\leq\Bigg\lceil\frac{\ln M+(d-2)\ln\!\Big(\frac{2eK}{d-2}\Big)}{\ln 2}\Bigg\rceil,$
which completes the proof. ∎
To bound $N_{ID}^{(2)}(\mathbf{I};K)$, we first relate it to $N_{ID}^{(2)}(M,I;K)$. The bound on the latter, provided by Theorem 3.30, then directly implies a bound on the former.
Corollary 3.31.
Let $\mathbf{I}\in\mathbb{N}_{0}^{M\times M}$, and let $I_{max}$ denote its largest entry. If $I_{max}\leq I$, then
$N_{ID}^{(2)}(\mathbf{I};K)\leq N_{ID}^{(2)}(M,I;K).$
4 VT-syndrome function
VT-syndrome function captures the weighted sum of bit positions for a given binary sequence and is used to define VT Code that is used to correct single insdel errors.
Definition 4.1 (VT-syndrome function).
A VT-syndrome function is defined as $f(u_{i})=VT(u_{i})=\sum_{j=1}^{k}ju_{ij}\pmod{k+1}$, $\forall$ $u_{i}=\{u_{i1},\ldots,u_{ik}\}\in\mathbb{F}^{k}_{2}$ and $\forall$ $i\in[2^{k}]$ such that $E=|Im(f)|=k+1.$
Using Theorem 3.22 and Corollary 3.19, we establish an upper bound and lower bound on $r^{VT}_{ID}(k,t)$ in the next lemma.
Lemma 4.2.
Let $f:\mathbb{F}_{2}^{k}\rightarrow\{0,1,\ldots,k\}$ be the VT-syndrome function defined by $f_{i}=i-1$ for all $1\leq i\leq k+1$, and let $x_{1}=(0,\ldots,0)$, $x_{i}=(0^{i-2}10^{k-i+1})$ denote the set of $k+1$ representative binary vectors of length $k$, where $2\leq i\leq k+1$. Consider the function distance matrix $\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)$ corresponding to $f$, and the distance matrix $I_{f}^{(1)}(t,x_{1},\ldots,x_{k+1})$ corresponding to the set of representative binary vectors. Then,
$N_{ID}^{(1)}(I_{\text{ID}}^{(1)}(t,x_{1},\ldots,x_{k+1}))\;\leq\;r^{\text{VT}}_{\text{ID}}(k,t)\;\leq\;N_{ID}^{(2)}(\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)),$
where
$[\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2(t+k),&i\neq j,\end{cases}$
and
$[I_{\text{ID}}^{(1)}(t,x_{1},\ldots,x_{k+1})]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2t,&i\neq j.\end{cases}$
Proof.
Consider the set of $k+1$ binary vectors $\{x_{i}\}_{i=1}^{k+1}$ defined by
$x_{i}=(0,0,\ldots,\underbrace{1}_{(i-1)\text{th position}},\ldots,0),\quad 1\leq i\leq k+1.$
The function values corresponding to these vectors are given by
$f_{i}=f(x_{i})=i-1,\quad\forall\,1\leq i\leq k+1.$
For any distinct pair $(x_{i},x_{j})$, the insertion–deletion distance satisfies
$d_{\text{ID}}(x_{i},x_{j})=2(k-(k-1))=2.$
Hence, the corresponding functional insdel distance satisfies
$2\leq d_{\text{ID}}^{f}(f_{i},f_{j})\leq d_{\text{ID}}(x_{i},x_{j})=2,\quad\forall\,1\leq i\neq j\leq k+1.$
Therefore, the entries of the function insdel matrix and insdel distance matrices associated with the set $\{x_{i}\}_{i=1}^{k+1}$ are given by
$\bm{I}_{f}^{(2)}((t,f_{1},\ldots,f_{k+1});k)]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2(t+k),&i\neq j,\end{cases}$
and
$I_{\text{ID}}^{(1)}(t,x_{1},\ldots,x_{k+1})]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2t,&i\neq j.\end{cases}$
This establishes the desired bounds on $r^{\text{VT}}_{\text{ID}}(k,t)$, completing the proof. ∎
Remark 4.3.
Since the non-diagonal entries of both the distance matrices $(I_{f}^{(1)}(t,x_{1},\ldots,x_{k+1})$ and $I_{\text{ID}}^{(2)}(t,x_{1},\ldots,x_{k+1})$ are equal to the same value $2t$ and $2(t+k)$ respectively, the bound on the optimal redundancy in Lemma 4.2 can be expressed as:
$N_{ID}^{(1)}(k+1,2t)\;\leq\;r^{\text{VT}}_{\text{ID}}(k,t)\;\leq\;N_{ID}^{(2)}(k+1,2(t+k);k))$
The next corollary gives an explicit bound on the optimal redundancy of FCIDCs for the VT syndrome function. The upper bound can be derived by using the result of the preceding remark together with Theorem 3.30, whereas the lower bound follows from the Remark 4.3 and Lemma 3.28.
Corollary 4.4.
Let $k\geq 2$ and $t\in\mathbb{N}$, then the optimal redundancy of FCIDCs for the VT syndrome function defined over $\mathbb{F}_{2}^{k}$ is bounded above as follows:
$\frac{2t(k+1)}{k+2}\leq r^{\text{VT}}_{\text{ID}}(k,t)\leq\Bigg\lceil\frac{\ln(k+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil.$
The following example illustrates Lemma 4.2 for particular values of $k$ and $t$.
Example 4.5.
Let $k=2$ and $t=1$. Then, from Lemma 4.2, the function distance matrix corresponding to the VT-syndrome function and the distance matrix corresponding to the binary vectors $x_{1}=(00)$, $x_{2}=(10)$, and $x_{3}=(01)$ are given respectively by
$\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&6&6\\[4.0pt] 6&0&6\\[4.0pt] 6&6&0\end{pmatrix}$,
and
$I_{\text{ID}}^{(1)}(1,x_{1},x_{2},x_{3})]_{ij}=\begin{pmatrix}0&2&2\\[4.0pt] 2&0&2\\[4.0pt] 2&2&0\end{pmatrix}$.
Consider the code $\mathcal{C}=\{000000,000111,110100\}$. It can be easily verified that the insdel distance between each pair of codewords in $\mathcal{C}$ is $6$. Hence,
$r^{\text{VT}}_{\text{ID}}(2,1)\leq N_{f}^{(2)}(\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2))\leq 6$
5 Number of Runs Function
A run in a binary sequence is a maximal block of consecutive equal bits. One can describe a binary string $x$ in terms of runs, and it is called run-length sequence, $R(x)=(r_{1},r_{2},\ldots,r_{m})$.
Definition 5.1.
A total number of runs function is defined as $f(x)=r(x)$, where $r(x)=|R(x)|$, $x\in\mathbb{F}^{k}_{2}$, and $k\in\mathbb{N}$.
Example 5.2.
Consider the vector $u=0100101$, then the run-length vector corresponding to vector $u$ is $[1,1,2,1,1,1]$ and hence the total number of runs equals the length of the run-length vector, i.e., $r(u)=6$.
The expressiveness of this function is given by $E=|Im(r(\cdot))|=k$. Using Corollary 3.19 and Theorem 3.22, we establish bounds on the optimal redundancy of FCIDCs for the function $r(\cdot)$, as stated in the following lemma.
Lemma 5.3.
Let $f(x)=r(x)$ be the number of runs function on $x\in\mathbb{F}_{2}^{k}$. Consider the set of $k$ binary vectors defined as $x_{i}=0^{k-i+1}(10)^{(i-1)/2}$ if $i$ is odd and $x_{i}=0^{k-i+1}(10)^{(i-2)/2}1$ if $i$ is even. Then the optimal redundancy of FCIDCs corresponding to $f$ is bounded as follows
$N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))\leq r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\bm{I}_{f}^{(2)}(t,f_{1},f_{2},\dots,f_{k});k)$
Where $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$ and $\bm{I}_{f}^{(2)}(t,f_{1},f_{2},\dots,f_{k});k)$ are order $k$ distance matrix and function distance matrix, respectively, and their entries are as follows:
[…]
and
[…]
Proof.
Consider the set of $k$ binary vectors $\{x_{i}\}_{i=1}^{k}$ defined by
[…]
The function values corresponding to these vectors are given by
$f_{i}=f(x_{i})=i,\quad\forall\,1\leq i\leq k.$
For any distinct pair $(x_{i},x_{j})$, the insertion–deletion distance satisfies
$d_{\text{ID}}(x_{i},x_{j})=\begin{cases}|i-j|,&\text{ if $|i-j|$ is even}\\ |i-j|+1,&\text{if $|i-j|$ is odd}\end{cases}$
and for any two binary vectors, say $x$ and $y$ of the same length, the insdel distance between the two vectors is bounded below by the modulus of the difference of the number of runs of the respective vectors, i.e., $|r(x)-r(y)|\leq d_{ID}(x,y)$.
Hence, the corresponding function distance satisfies
$|i-j|\leq d_{\text{ID}}^{f}(f_{i},f_{j})\leq d_{\text{ID}}(x_{i},x_{j}),\quad\forall\,1\leq i\neq j\leq k.$
Since the insdel distance is always even, from the above inequality, we conclude
$d_{\text{ID}}^{f}(f_{i},f_{j})=d_{\text{ID}}(x_{i},x_{j})\quad\forall\,1\leq i\neq j\leq k$
Therefore, the entries of the function insdel matrix and insdel distance matrices associated with the set $\{x_{i}\}_{i=1}^{k}$ are given respectively by
[…]
and
[…]
This establishes the desired bounds on $r^{f}_{\text{ID}}(k,t)$, completing the proof. ∎
Example 5.4.
Let $k=4$ and $t=1$. Then, from Lemma 5.3, the function distance matrix corresponding to the number of runs function and the distance matrix corresponding to the binary vectors $x_{1}=(0000)$, $x_{2}=(0001)$, $x_{3}=(0010)$ and $x_{4}=(0101)$ are given respectively by
$\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&10&10&8\\[4.0pt] 10&0&10&10\\[4.0pt] 10&10&0&10\\[4.0pt] 8&10&10&0\end{pmatrix}$,
and
$I_{\text{ID}}^{(1)}(1,x_{1},x_{2},x_{3})]_{ij}=\begin{pmatrix}0&2&2&0\\[4.0pt] 2&0&2&2\\[4.0pt] 2&2&0&2\\[4.0pt] 2&2&2&0\end{pmatrix}$.
The next result presents a lower bound on the redundancy of the number of runs function that is inferred from Lemma 5.3 and Plotkin-like bound of Lemma 3.26
Corollary 5.5.
For any $k\geq t+2$. Then,
$r_{ID}^{f}(k,t)\geq\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right].$
This simplifies to:
$r_{\text{ID}}^{f}(k,t)\geq\frac{10t^{3}+39t^{2}+26t}{3(t+2)^{2}}\quad,\text{ when $t$ is even }$
and
$r_{\text{ID}}^{f}(k,t)\geq\frac{10t^{3}+39t^{2}+38t+29}{3(t+2)^{2}},\quad\text{ when $t$ is odd. }$
Proof.
For $1\leq i\leq k$, let $x_{i}=0^{k-i+1(10)^{(i-1)/2}}$ if $i$ is odd and $x_{i}=0^{k-i+1}(10)^{(i-2)/2}1$ if $i$ is even. Let $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$ be a $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$-code of length $N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))$. Consider the first $t+2$ codewords of $\mathcal{P}$ and $(t+2)\times(t+2)$ leading principal submatrix of $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k})$ denoted by $I_{t+2}.$ Then, $\{p_{1},p_{2},\ldots,p_{t+2}\}$ is a $I_{t+2}$-code. By Corollary 3.27,
$N_{ID}^{(1)}[I_{t+2}]$ $\geq\frac{2}{(t+2)^{2}}\sum_{i=1}^{t+2}\sum_{j=i+1}^{t+2}[I_{t+2}]_{ij}$
$=\frac{2}{(t+2)^{2}}\left(\sum_{\begin{subarray}{c}i=0\\ i\text{ is even}\end{subarray}}^{t}(t+1-i)(2t-i)+\sum_{\begin{subarray}{c}i=0\\ i\text{ is odd}\end{subarray}}^{t}(t+1-i)(2t+1-i)\right)$
$=\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right]$
From Lemma 5.3, we have
$r_{ID}^{f}(k,t)$ $\geq N^{(1)}_{ID}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\dots,x_{k}))$
$\geq N_{ID}^{(1)}[I_{t+2}]$
$\geq\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right]$
The bounds for the cases of odd and even $t$ are immediate consequences of the preceding bound. ∎
Next, we present a construction for FCIDCs for the number of runs function which utilises the concept of the shifted modulo operator defined in [1].
The shifted modulo operator is defined as follows:
Definition 5.6.
$a\text{ }smod\text{ }b\triangleq((a-1)\pmod{b})+1\in\{1,2,\ldots,b\}$
Construction 1.
Define
$\psi(x)=(x,p_{r(x)}),$
where the $p_{i}$’s are defined depending on $t$ as follows.
For $t=1$, set $p_{1}=(00)^{k+1},\text{ }p_{2}=(10)^{k+1},$ and $p_{3}=(0^{k+1}1^{k+1})$. Set $p_{i}=p_{i\text{ }smod\text{ }3}$ for $i\geq 4$.
For $t=2$, set $p_{1}=(00)^{k+2},\text{ }p_{2}=(10)^{k+2},$ $p_{3}=(0^{k+2}1^{k+2})$ and $p_{4}=(1^{k+2}0^{k+2})$. Set $p_{i}=p_{i\text{ }smod\text{ }5}$ for $i\geq 5$.
For $t\geq 3$, let $p_{1},\ldots,p_{2t+1}$ be an insdel code with minimum insdel distance $2(k+t)$, i.e., $d_{ID}(p_{i},p_{j})\geq 2(t+k)\quad\forall i,j\leq 2t+1,\text{ }i\neq j$ and set $p_{i}=p_{i\text{ }smod\text{ }(2t+1)}$ for $i\geq 2t+2$.
The FCIDC construction gives an upper bound on the optimal redundancy of the number of runs function as stated in the following lemma.
Lemma 5.7.
Let $k,t\in\mathbb{N}$ and $f$ be the number of runs function defined on $\mathbb{F}_{2}^{k}$. Then, the optimal redundancy of FCIDCs corresponding to f is upper bounded as follows:
  • $r_{ID}^{f}(k,1)\leq 2(k+1)$.
  • $r_{ID}^{f}(k,2)\leq 2(k+2)$.
  • $r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(2t+1,2(t+k);k)\leq\Bigg\lceil\frac{\ln(2t+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$ .
The proof of this lemma follows from the construction given in the Appendix 9.
6 Maximum Run-Length Function
Another interesting function that can be defined from the run-length sequence $R(x)$ is the maximum run-length function , which gives the length of the longest block of consecutive identical bits in a given vector.
Definition 6.1 (Maximum Run-Length Function).
Let $x\in\mathbb{F}_{2}^{k}$ and $R(x)=(r_{1},r_{2},\ldots,r_{m})$ then the maximum run-length function is defined as:
$r_{max}(x)=max_{1\leq i\leq m}r_{i}$
Example 6.2.
Consider the vector $u=0000101$ , then the run-length vector corresponding to the vector $u$ is $(4,1,1,1)$ , and hence the maximum run-length of $u$ is $4.$
The expressiveness for this function is given by $E=|Im(r_{max}(\cdot))|=|\{1,2,\dots,k\}|=k$.
The following result establishes a lower bound on the function distance for the maximum run-length function.
Lemma 6.3.
Let $f:\mathbb{F}_{2}^{k}\rightarrow\{1,2,\ldots,k\}$ be the maximum run-length function. Then,
$d_{ID}^{f}(i,j)\geq 2\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\text{ for all }1\leq i,j\leq k.$
Proof.
Let $x$ be a binary vector of length $k$ and maximum run-length $i$. Without loss of generality, we assume that the longest run of $x$ is $0^{i}$. To derive the lower bound, we consider a binary vector $y$ of length $k$ and maximum run-length $j$ that is as close as possible to $x$ while respecting the maximum run-length constraint. In this way, the minimal possible insdel distance will occur when the long runs in $x$ and $y$ are of the same symbol.
Let $y^{\prime}$ be a substring of $y$ aligned with the run $0^{i}$. If $t$ is the number of $1s$ in $y^{\prime}$, then the number of zeros in $y^{\prime}$ is iti-t. This means $y^{\prime}$ can have at most $t+1$ blocks of zeros, each split by $1$ and having size at most $j$. Hence, we have,
$i-t\leq(t+1)j\implies t\geq\Big\lceil\frac{i-j}{j+1}\Big\rceil$
which means in order to have a run of $0s$ of length $j$, $y^{\prime}$ should have at least $\Big\lceil\frac{i-j}{j+1}\Big\rceil$ number of $1$. Hence,
$LCS(x^{\prime},y^{\prime})\geq i-\Big\lceil\frac{i-j}{j+1}\Big\rceil$
as $LCS(x^{\prime},y^{\prime})$ is equal to the maximum number of $0s$ in $y^{\prime}$. The length of longest common subsequence of $x$ and $y$ can be bounded above in terms of $LCS(x^{\prime},y^{\prime})$ as follows:
$LCS(x,y)\leq LCS(x^{\prime},y^{\prime})+(k-i)$
From the formula of the insdel distance in terms of $LCS$ we get:
$d_{ID}(x,y)$ $=2k-2LCS(x,y)$
$\geq 2k-2(LCS(x^{\prime},y^{\prime})+(k-i))$
$\geq 2\Big\lceil\frac{i-j}{j+1}\Big\rceil.$
Remark 6.4.
The above bound is tight; that is, for any integers $i,j$ with $0<j<i\leq k$, there exist binary vectors $x$ and $y$ of length $k$ with $r_{\max}(x)=i$ and $r_{\max}(y)=j$ for which equality is achieved.
The following example shows that the bound in Lemma 6.3 is tight.
Example 6.5.
Let $k=5$ and $\mathcal{S}=\{00000,00001,00010,00100,01010\}$. Then for any pair of strings $x,y$ in $\mathcal{S}$ having maximum run-length $i$ and $j$ respectively, $d_{ID}(x,y)=2\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\text{ for all }1\leq i,j\leq 5.$
Using Lemma 6.3 and Corollary 3.24, we give an upper bound on the optimal redundancy of FCIDCs designed for maximum run-length function.
Lemma 6.6.
Let $f(x)=r_{max}(x)$ defined over $\mathbb{F}_{2}^{k}$. Then,
$r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(\bm{I};k).$
where
$\bm{I}]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2\Big(t+k+1-\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\Big),&i\neq j\end{cases}$
Proof.
By Lemma 6.3, $d_{ID}(x,y)\leq 2\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\quad\forall 1\leq i,j\leq k.$ Let $\bm{I}$ be a square matrix of order $k$ whose entries are:
$\bm{I}]_{ij}=\begin{cases}0,&i=j,\\[4.0pt] 2\Big(t+k+1-\Big\lceil\frac{|i-j|}{min(i,j)+1}\Big\rceil\Big),&i\neq j\end{cases}$
Then, by Corollary 3.24 $r_{ID}^{f}(k,t)\leq N_{ID}^{(2)}(\bm{I};k).$
Observe that $max_{i,j}[\mathbf{I}]_{ij}=2(t+k)$ in Lemma 6.6, therefore the following redundancy bound follows from Corollary 3.31.
Corollary 6.7.
Let $k,t\in\mathbb{N}$ and $r_{ID}^{f}(k,t)$ be the optimal redundancy of FCIDCs corresponding to $f(x)=r_{max}(x)$. Then,
$N_{ID}^{(2)}(k,2(t+k);k)\leq\Bigg\lceil\frac{\ln(k)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil.$
The next lemma gives a lower bound on the optimal redundancy of FCIDCs designed for the maximum run-length functions by using Corollary 3.19.
Lemma 6.8.
Let $f(x)=r_{max}(x)$ and $x_{i}=0^{i}(10)^{(k-i)/2}$ if kik-i is even and $x_{i}=0^{i}(10)^{(k-i-1)/2}1$ if kik-i is odd for $i\in\{1,2,\ldots,k\}$.Then,
$r_{ID}^{f}(k,t)\geq N_{ID}^{(1)}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\ldots,x_{k})).$
where
[…]
Proof.
Consider the following set of representative vectors: $x_{i}=0^{i}(10)^{(k-i)/2}$ if kik-i is even and $x_{i}=0^{i}(10)^{(k-i-1)/2}1$ if kik-i is odd for $i\in\{1,2,\ldots,k\}$. Then, $f(x_{i})=r_{max}(x_{i})=i$. We claim that the insdel distance between these representative set of vectors is given by:
$d_{\text{ID}}(x_{i},x_{j})=\begin{cases}|i-j|,&\text{ if $|i-j|$ is even}\\ |i-j|+1,&\text{if $|i-j|$ is odd}\end{cases}$
WLOG assume $k$ is even and $i>j$.
Case 1: Both $i$ and $j$ are even.
Then, $\operatorname{LCS}(x_{i},x_{j})\geq j+\frac{i-j}{2}+k-i=k-\frac{i-j}{2}.$
Hence, $d_{ID}(x_{i},x_{j})=2(k-\operatorname{LCS}(x_{i},x_{j}))\leq i-j$.
From Lemma 2.9 we have $d_{ID}(x_{i},x_{j})\geq|r(x_{i})-r(x_{j})|=|(1+k-i)-(1+k-j)|=|j-i|=i-j.$
Hence, $d_{ID}(x_{i},x_{j})=i-j$.
Case 2: Both $i$ and $j$ are odd.
Then, $\operatorname{LCS}(x_{i},x_{j})\geq j+\frac{i-j}{2}+(k-i-1)+1=k-\frac{i-j}{2}.$
Therefore, using the same argument as in Case 1, we get $d_{ID}(x_{i},x_{j})=i-j$.
Case 3: $i$ is even $j$ is odd.
Then, $\operatorname{LCS}(x_{i},x_{j})\geq j+\frac{i-j-1}{2}-1+(k-i-1)=k-\frac{i-j-1}{2}+1.$
Hence, $d_{ID}(x_{i},x_{j})=2(k-\operatorname{LCS}(x_{i},x_{j}))\leq i-j+1$.
From Lemma 2.9 we have $d_{ID}(x_{i},x_{j})\geq|r(x_{i})-r(x_{j})|+1=|(1+k-i)-(1+k-j)|+1=|j-i|+1=i-j+1.$
Therefore, using the same argument as in Case 1, we get $d_{ID}(x_{i},x_{j})=i-j+1$.
Case 4: $i$ is odd $j$ is even.
Using the same argument of Case 3, one can easily verify that $d_{ID}(x_{i},x_{j})=i-j+1$.
Therefore, we can conclude that
[…]
and the lower bound on the redundancy, $r_{ID}^{f}(k,t)\geq N_{ID}^{(1)}(\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\ldots,x_{k}))$ follows from Corollary 3.19.
Because the matrix entries $\bm{I}_{f}^{(1)}(t,x_{1},x_{2},\ldots,x_{k})$(Lemma 6.8) associated with maximum run length function match those for the number of runs (Lemma 5.3), the lower bound on optimal redundancy for FCIDCs is identical in both cases(Corollary 5.5), as given in the corollary below.
Corollary 6.9.
For any $k\geq t+2$. Then,
$r_{ID}^{f}(k,t)\geq\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right].$
This simplifies to:
[…]
and
[…]
Example 6.10.
Let $k=4$ and $t=1$. Then, from Lemma 5.3, the function distance matrix corresponding to the number of runs function and the distance matrix corresponding to the binary vectors $x_{1}=(0000)$, $x_{2}=(0001)$, $x_{3}=(0010)$ and $x_{4}=(0101)$ are given respectively by
$\bm{I}_{f}^{(2)}((1,f_{1},f_{2},f_{3});2)]_{ij}=\begin{pmatrix}0&10&10&8\\[4.0pt] 10&0&10&10\\[4.0pt] 10&10&0&10\\[4.0pt] 8&10&10&0\end{pmatrix}$,
and
$I_{\text{ID}}^{(1)}(1,x_{1},x_{2},x_{3})]_{ij}=\begin{pmatrix}0&2&2&0\\[4.0pt] 2&0&2&2\\[4.0pt] 2&2&0&2\\[4.0pt] 2&2&2&0\end{pmatrix}$.
7 Locally $(\lambda,\rho)_{ID}$-function
Lenz et al. first studied a class of functions, called locally binary function[1] with respect to the Hamming metric and derived optimal function-correcting codes (FCCs) for them. This concept was subsequently generalized by Rajput et al. to locally $(\lambda,\rho)$-functions[16] for the Hamming metric, and later by Verma et al. for the $b$-symbol metric [18]. The significance of this class of functions stems from the fact that any function with a finite image set can be represented as a locally $(\lambda,\rho)$-function. This universality is the primary reason the theory, initially developed for the Hamming metric, has been extended to other distance metrics where FCCs are being explored. In this section, we study the class of locally $(\lambda,\rho)_{ID}$-functions in the insdel metric setting.
Definition 7.1 (Function ball).
The function ball of a function $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ with radius $\rho$ around $u\in\mathbb{F}_{2}^{k}$ in insdel metric is defined as
$B_{ID}^{f}(u,\rho)=f(B_{ID}(u,\rho))=\{f(v)|v\in B_{ID}(u,\rho)\}.$
Definition 7.2 (Locally bounded function in insdel metric).
A locally $(\lambda,\rho)_{ID}$ function $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ is a function for which $|B_{ID}^{f}(u,\rho)|\leq\lambda,\quad\forall u\in\mathbb{F}_{2}^{k}.$
The following lemma is a straightforward extension of Lemma 2.17 to the insdel metric setting. This result has subsequently been employed to derive an upper bound on the optimal redundancy of FCIDCs corresponding to locally $(\lambda,\rho)_{ID}$-functions.
Lemma 7.3.
Let $f:\mathbb{F}_{2}^{k}\rightarrow\mathrm{Im}(f)$ be a locally $(\lambda,\rho)_{ID}$-function. Assume that $\mathrm{Im}(f)$ is equipped with a total order $\prec$, and that for every $u\in\mathbb{F}_{2}^{k}$, the set $B_{ID}^{f}(u,\rho)$ forms a contiguous block with respect to $\prec$. Then there exists a mapping
$\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[\lambda]$
such that for all $u,v\in\mathbb{F}_{2}^{k}$ satisfying $f(u)\neq f(v)\quad\text{and}\quad d_{ID}(u,v)\leq\rho$, we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$.
Throughout the remainder of this section, we assume that all locally $(\lambda,\rho)_{ID}$ functions under consideration satisfy the assumption in Lemma 7.3. We now present an upper bound on the optimal redundancy of FCIDCs corresponding to locally $(\lambda,\rho)_{ID}$-functions, expressed in terms of the shortest possible length of a binary insdel code with a prescribed number of codewords and minimum insdel distance.
Theorem 7.4.
Let $t$ be a positive integer and $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a locally $(\lambda,2t)_{ID}$-function then $r_{ID}^{f}(k,t)\leq N^{(2)}_{ID}(\lambda,2(t+k);k)\leq\frac{\ln\lambda+2(t+k-1)\ln\frac{ek}{t+k-1}}{\ln 2}$.
Proof.
Let $f$ be a locally $(\lambda,2t)_{ID}$-function. Then by Lemma 7.3, there exists a mapping
$\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\to[\lambda]$
such that for any $u,v\in\mathbb{F}_{2}^{k}$ satisfying $f(u)\neq f(v)$ and $d_{ID}(u,v)\leq 2t$, we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$. Let $\mathcal{C}$ be a binary insdel code of size $\lambda$ , minimum distance $2(t+k)$, and length $N_{ID}^{(2)}((\lambda,2(t+k));k)$. Denote the codewords of $\mathcal{C}$ as $C_{1},C_{2},\ldots,C_{\lambda}$ and define the encoding function
$\psi:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{k+N_{ID}^{(2)}(\lambda,2(t+k);k)}$
by
$\mathrm{\psi}(u)=(u,u_{p}),\quad\text{where }u_{p}=C_{\mathrm{Col}_{f}(u)}.$
We now show that the above encoding function defines an FCIDC with redundancy $r=N_{ID}^{(2)}((\lambda,2t);k)$. Let $u,v\in\mathbb{F}_{2}^{k}$ be such that $f(u)\neq f(v)$. There are two possible cases for $u$ and $v$.
Case 1: If $d_{ID}(u,v)>2t$, then
$d_{ID}(\mathrm{\psi}(u),\mathrm{\psi}(v))\geq d_{ID}(u_{p},v_{p})>2t.$
Case 2: If $d_{ID}(u,v)\leq 2t$, then by the definition of $\mathrm{Col}_{f}$ we have $\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)$. Therefore,
$d_{ID}(u_{p},v_{p})=d_{ID}(C_{\mathrm{Col}_{f}(u)},C_{\mathrm{Col}_{f}(v)})\geq 2(t+k),$
Because $u\neq v$, we have $d_{ID}(u,v)\geq 2$, and thus
$d_{ID}(\mathrm{\psi}(u),\mathrm{\psi}(v))$ $\geq d_{ID}(u,v)+d_{ID}(u_{p},v_{p})-2\min{(k,r)}$
$\geq 2+2(t+k)-2k$
$=2t+2.$
The second inequality in the above equation follows from the fact that $N_{ID}^{(2)}((\lambda,2(t+k));k)=r>k$ .
Hence, in both cases, the encoding satisfies the desired distance property, proving that $\psi$ defines an FCIDC with redundancy $r=N_{ID}^{(2)}((\lambda,2t);k)$. ∎
Corollary 7.5.
Let $t$ be a positive integer and $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a $2t$-local binary insdel function then
$t\leq r_{ID}^{f}(k,t)\leq t+k.$
Proof.
The lower bound follows from Corollary 3.19, while the upper bound is obtained from Theorem 7.4 together with the observation that $N^{(2)}_{ID}(2,2(t+k);k)=t+k$. ∎
Corollary 7.6.
Let $t$ be a positive integer and $f:\mathbb{F}_{2}^{k}\rightarrow Im(f)$ be a $(2t,3)$-locally insdel function then
$r_{ID}^{f}(k,t)\leq 2(t+k).$
Proof.
From Lemma 7.3, there exists a mapping $\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[3]$ corresponding to the function $f$, such that for any $u,v\in\mathbb{F}_{2}^{k}$,
$\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)\quad\text{whenever}\quad f(u)\neq f(v)\ \text{and}\ d_{ID}(u,v)\leq 2t.$
Consider the following encoding function
$\psi:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+2(k+t)}$
defined as follows: $\psi(u)=(u,u_{p})$ where $u_{p}=(u_{p}^{\prime})^{t+k}$ and
$u_{p}^{\prime}=\begin{cases}00\quad if\quad Col_{f}(u)=1\\ 01\quad if\quad Col_{f}(u)=2\\ 11\quad if\quad Col_{f}(u)=3.\end{cases}$
It can easily verified that $LCS((00)^{t+k},(01)^{t+k})=t+k$ as both the vectors have $t+k$ common $0s$. Similarly, $LCS((11)^{t+k},(01)^{t+k})=t+k$ and $LCS((00)^{t+k},(11)^{t+k})=0$. Therefore, the insdel distance between any distinct redundancy vector is at least $2(t+k)$. We now prove that the above encoding function $\psi$ is an FCIDC with redundancy $2(t+k)$.
Let $u,v\in\mathbb{F}_{2}^{k}$ be such that $f(u)\neq f(v).$ Consider the following two cases with respect to the insdel distance between $u$ and $v.$
Case 1: $d_{ID}(u,v)>2t$
$d_{ID}(\psi(u),\psi(v))\geq d_{ID}(u,v)>2t$
Case 2: $d_{ID}(u,v)\leq 2t$
$d_{I}(\psi(u),\psi(v))$ $\geq d_{I}(u,v)+d_{I}(u_{p},v_{p})-2\cdot min(k,2(k+t))$
$\geq 2+2(t+k)-2k$
$\geq 2t+2.$
Given that any function having a finite image set can be expressed as a locally $(\lambda,\rho)_{ID}$-function, we next analyze the number of runs functions as locally $(2t,\lambda)$-bounded functions.
Proposition 7.7.
Let $t\in\mathbb{N}$ then the number of runs function is a locally $(2t,4t+1)_{ID}$-function, i.e.
$|B^{f}_{ID}(x,2t)|\leq 4t+1,\quad\forall x\in\mathbb{F}_{2}^{k}$
Proof.
Let $f=r(x)$ be a number of runs function. We claim that
$|B^{f}_{ID}(x,2t)|\leq 4t+1.$
Suppose, for the sake of contradiction, that there exists some $x\in\mathbb{F}_{2}^{k}$ such that
$|B_{ID}^{f}(x,2t)|>4t+1.$
Then we can find two vectors $y_{1},y_{2}\in\mathbb{F}_{2}^{k}$ with $d_{ID}(x,y_{1})\leq 2t$, $d_{ID}(x,y_{2})\leq 2t$, and
$f(y_{1})=\max B_{ID}^{f}(x,2t),\qquad f(y_{2})=\min B_{ID}^{f}(x,2t).$
Clearly,
$f(y_{1})-f(y_{2})\geq 4t+1.$ (7.1)
On the other hand, using the results of Lemma 2.9 and the triangle inequality, we get,
${r}(y_{1})-r(y_{2})\leq d_{ID}(y_{1},y_{2})\leq d_{ID}(x,y_{1})+d_{ID}(x,y_{2})\leq 4t.$ (7.2)
Thus from equation 7.1 and 7.2 we have,
$4t\geq f(y_{1})-f(y_{2})\geq 4t+1.$
which is a contradiction.
Therefore,
$|B_{ID}^{f}(x,2t)|\leq 4t+1\qquad\text{for all }x\in\mathbb{F}_{2}^{k}.$
Hence, the function $f$ is locally $(2t,4t+1)_{ID}$-function. ∎
Function Lower Bound Upper Bound
VT-syndrome function $\frac{2t(k+1)}{k+2}$ $\Bigg\lceil\frac{\ln(k+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$.
Number of runs function $\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right]$ $\Bigg\lceil\frac{\ln(2t+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$
Maximum run-length function $\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right]$ $\Bigg\lceil\frac{\ln(k)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$
Locally binary insdel function $t$ $t+k$
Table 1: Lower bound and upper bound on the optimal redundancy of FCIDCs for different functions.
8 Conclusion
This work introduces a comprehensive and unified framework for function-correcting codes designed specifically for insertion–deletion (insdel) channels. We develop a general construction methodology for such codes and demonstrate its applicability through several representative classes of functions, supported by concrete examples. To characterize the fundamental limits of these codes, we establish Gilbert–Varshamov and Plotkin-like bounds on the lengths of irregular insdel-distance codes. By exploiting the intrinsic relationship between the optimal redundancy of function-correcting insdel codes and the minimal lengths of these irregular distance codes, we further derive a simplified and broadly applicable lower bound on redundancy. Building on these theoretical insights, we explicitly construct FCIDCs for various function families, including locally bounded functions, VT-type functions, the number-of-runs function, and the maximum-run-length function, thereby demonstrating the versatility and practical potential of our framework. Finally, we highlight that extending function-error correction to DNA data storage, where insertion and deletion errors are particularly prevalent, offers a promising direction for future research.
Acknowledgment
We are grateful to Prof. Eitan Yaakobi for his valuable discussions and suggestions, which helped improve the current version of the manuscript.
9 Appendix
We show that the construction 1 yields an FCIDC for the redundancy vectors.
For $t=1$, assume $r(x_{i})\neq r(x_{j})$ and consider the following two cases.
Case 1: $r(x_{i})\equiv r(x_{j})\text{ }smod{3}$.
Since the run counts belong to the same residue class modulo 3 and are distinct, the smallest possible run count difference is 3.
From Lemma 2.9 we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)\geq 2\Bigl\lceil\frac{3}{2}\Bigr\rceil=4.$
Thus, whenever $r(x_{i})\equiv r(x_{j})\text{ }smod{3}$ and $r(x_{i})\neq r(x_{j})$,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 4.$
Case 2: $r(x_{i})\not\equiv r(x_{j})\text{ }smod{3}$.
Here the redundancy parts $p_{r(x_{i})}$ and $p_{r(x_{j})}$ are distinct codewords code with minimum insdel distance $2(k+1)$. Hence we have,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)$ $=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)$
$\geq d_{ID}(x_{i},x_{j})+d_{ID}(p_{r(x_{i})},p_{r(x_{j})})-2\min\{k,2(k+1)\},$
Since, $d_{ID}(x_{i},x_{j})\geq 2$, $d_{ID}(p_{r(x_{i})},p_{r(x_{j})})\geq 2(k+1)$ and $\min\{k,2(k+1)\}=k$, we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 2+2(k+1)-2k=4.$
For $t=2$ a completely analogous analysis shows that the concatenation of message vectors with the corresponding redundancy vectors yields a FCIDC.
For $t\geq 3$, assume $r(x_{i})\neq r(x_{j})$ and consider the following two cases.
Case 1: $r(x_{i})\equiv r(x_{j})\text{ }smod{(2t+1)}$.
Since the run counts belong to the same residue class smodsmod $2t+1$ and are distinct, the smallest possible run count difference is $2t+1$.
From Lemma 2.9 we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)\geq 2\Bigl\lceil\frac{2t+1}{2}\Bigr\rceil=2t+2.$
Thus, whenever $r(x_{i})\equiv r(x_{j})\text{ }smod{(2t+1)}$ and $r(x_{i})\neq r(x_{j})$,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 2t+2.$
Case 2: $r(x_{i})\not\equiv r(x_{j})\text{ }smod{(2t+1)}$.
Here the redundancy parts $p_{r(x_{i})}$ and $p_{r(x_{j})}$ are distinct codewords code with minimum insdel distance $2(k+t)$. Hence we have,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)$ $=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)$
$\geq d_{ID}(x_{i},x_{j})+d_{ID}(p_{r(x_{i})},p_{r(x_{j})})-2\min\{k,2(k+1)\},$
Since, $d_{ID}(x_{i},x_{j})\geq 2$, $d_{ID}(p_{r(x_{i})},p_{r(x_{j})})\geq 2(k+1)$ and $\min\{k,2(k+1)\}=k$, we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 2+2(k+t)-2k=2t+2.$